I have a question and hope it does not seem like a silly one. I am studying quadratic equations and the different type of ways to solve them. My question is:
Since the quadratic formula when used to solve quadratic equations always spits out the correct answer, why do we not use only this method rather than even bothering with the other methods?

Answer by algebrahouse.com(1659)

Since the quadratic formula when used to solve quadratic equations always spits out the correct answer, why do we not use only this method rather than even bothering with the other methods?

It requires less steps using other methods. Therefore, being more efficient.

For example:
x² = 49 is a quadratic equation
x = ±7 is the solution {achieved by simply taking the square root of both sides}

x² - 6x + 8 = 0 is a quadratic equation
(x - 4)(x - 2) = 0 {factored into two binomials}
x - 4 = 0 or x - 2 = 0 {set each factor equal to 0}
x = 4 or x = 2 {solved each equation for x}
Once again, less steps than the quadratic formula!
www.algebrahouse.com